# Neural Concept Binder

### v1

@@ 4.1 Evaluations
Large, gray cube C3 $\land$ H4 $\land$ K4 $\land$ O12 $\land$ P5 (Gray1) $\land$ (Red $\lor$ Gray2) $\land$ (Large) $\land$ (Gray3) $\land$ (Gray4) “A large gray object”
Small, metal cube B3 $\land$ D3 $\land$ H0 $\land$ I0 $\land$ K0 (Cube) $\land$ (Small1) $\land$ (Small2) $\land$ (Small3) $\land$ (Small4) “A small cube
Large, blue sphere B0 $\land$ C6 $\land$ H3 $\land$ O0 $\land$ P1 (Sphere) $\land$ (Blue1) $\land$ (Blue2) $\land$ (Small $\lor$ Blue3) $\land$ (Blue4 $\lor$ Green $\lor$ Purple) “A blue sphere
Utilising unsupervised concepts for understanding neural computations (Q4). In our final evaluations, we investigate whether NCB’s concept encodings can make subsymbolic computations more transparent. Therefore, we consider the task of image classification on variations of the benchmark CLEVR-Hans3 dataset [59]. While the concept encodings in NCB + NN are trained unsupervised, they perform on par with those of the supervised approach of [59] (cf. Sec. F.7). Even more important, integrating NCB’s inherently inspectable concept representations into neural computations can ellicit more transparent decision processes. We illustrate this in Tab.  3, where we exemplify class-level
@@
explanations of the classifier of NCB + NN (cf. Sec. E.4 for details). Via NCB’s inspection mechanisms, a human stakeholder can now easily identify the classifier’s internal decision rule for a class (e.g., "a large gray object"). Overall, this is a vital feature for the deployment of trustworthy AI models in real-world scenarios and the significance of our results is, that we can achieve this even when utilising unsupervised concept encodings. Moreover, in Fig.  6 we investigate whether a NCB-based neural classifier can be revised to mitigate the confounders in CLEVR-Hans3 (cf. Sec. E.4 and Sec. F.8 for details). The confounding factor in the training set is the color gray. We present the non-confounded test set accuracy in Fig.  6. We observe that standard loss-based feedback via explanatory interactive learning (XIL) [59] on the NN classifier’s explanations (+ XIL on NN) can greatly reduce the effect of the confounder. Alternatively, by simply zeroing the activations of the undesired concept gray (+ XIL on concepts) we even achieve improved confounding mitigation results without the usual issues of joint optimization. Overall, our results highlight the potential of integrating NCB’s unsupervised concept representations for eliciting transparent and trustworthy subsymbolic computations. We thus answer Q4 affirmatively.
Limitations. NCB largely benefits from qualitative initial block-slot encodings. If these are lacking, the extracted concept-slot encodings lose quality as well. An important next step to handle more complex visual inputs, e.g., video data, is the integration of recent approaches, e.g., [11, 14]. Furthermore, due to the unsupervised nature of NCB’s training, additional alignment of NCB’s concepts is inevitable for deployment in downstream tasks [5]. Further, to build trust in NCB’s concept knowledge, human inspection is required. Lastly, revisions are an integral part of NCB. However they rely on humans to provide "good" feedback, as a malicious user can manipulate NCB’s concepts. However, inspecting the concept space allows to track this effectively.
@@ 5 Conclusions
In this work, we introduce the Neural Concept Binder framework for learning visual object-factor concepts in an unsupervised fashion. Our evaluations suggest that NCB’s specific binding mechanisms benefit learning expressive yet discrete concept representations. Moreover, our evaluations highlight the potential of integrating NCB’s inherently inspectable and revisable concept-slot encodings into symbolic and neural modules. Interesting avenues for future research are investigating the benefits of NCB’s concept representations in the context of continual learning settings [7] as well as integrating NCB’s basic level concepts into high-level concept learning approaches [71], but also probabilistic logic programming approaches [57, 58]. Additionally, integrating learning signals from a downstream task into NCB’s concept representations, e.g., via differentiable clustering [67], is important for improving the quality of initial concept encodings.
@@ Acknowledgments
The authors wish to thank Gautam Singh for initial help with SysBinder. Furthermore, the authors wish to thank Cyprien Dzialo for his preliminary results and insights on this project. This work was supported by the Priority Program (SPP) 2422 in the subproject “Optimization of active surface design of high-speed progressive tools using machine and deep learning algorithms“ funded by the German Research Foundation (DFG), the ”ML2MT” project from the Volkswagen Stiftung and the ”The Adaptive Mind” project from the Hessian Ministry of Science and Arts (HMWK). It has further benefited from the HMWK projects ”The Third Wave of Artificial Intelligence - 3AI”, and Hessian.AI, as well as the Hessian research priority program LOEWE within the project WhiteBox, the ICT-48 Network of AI Research Excellence Center “TAILOR” (EU Horizon 2020, GA No 952215) and the EU-funded “TANGO” project (EU Horizon 2023, GA No 57100431).
@@ Appendix A Details on Neural Concept Binder
A.1 Selection Function
@@ A.2 Details on Training
A.3 Instantiating Neural Concept Binder
We instantiate NCB’s soft binder via the SysBinder approach of  Singh et al. [56] which has been shown to provide valuable, object-factor disentangled representations. Thus, the soft binder was trained as in the original setup and with the published hyperparameters. Furthermore we instantiate the clustering model, $h$, via the powerful HDBSCAN method [8, 9, 45] (based on the popular HDBSCAN library [^3] https://hdbscan.readthedocs.io/en/latest/index.html). Hereby, $h$’s internal representation, $\phi$, consists of the learned hierarchical merge tree. In practice we found it beneficial to perform a grid search over $h$’s hyperparameters based on the unsupervised density-based cluster validity score [44]. The searched parameters are the minimal cluster size (the minimum number of samples in a group for that group to be considered a cluster) and minimal sample number (the number of samples in a neighborhood for a point to be considered as a core point) each over the values $[5,10,15,20,25,30,50,80,100]$. Moreover, we utilize the excess of mass algorithm and allow for single clusters. We performed the training of the retrieval corpus, i.e., fitting $h$, on a dataset of images containing single objects for simplifying the subsequent concept inspection mechanisms of our evaluations. However, this can easily be extended to multiple object images by utilising the soft binder’s slot attention masks to identify relevant objects in an image. Finally, we instantiate the retrieval corpus as a set of dictionaries and, unless stated otherwise, we utilise a retrieval corpus which contains one prototype and a set of exemplar encodings per concept. Furthermore, $s_{\mathcal{R}}$ represents the $\operatorname*{argmin}$ selection function and we utilize the euclidean distance as $d(\cdot,\cdot)$. It is important to note that $h$ does not make any assumptions about the number of clusters, $N_{C}$. Thus, although $h$ fits a clustering to best fit the block-slot encodings of a block, it can potentially provide an overparameterized clustering, e.g. by representing one underlying factor such as “gray” with several clusters. This highlights the importance of task-alignment, e.g., for symbolic downstream tasks, and concept inspection for general concept alignment. We refer to our code for more details [^4] Code available here., where we will publish trained model checkpoints and corresponding parameter logs upon acceptance.
A.4 Computational Resources
@@ Appendix B Details on CLEVR-Sudoku
CLEVR-Sudoku provides Sudokus based on the datasets CLEVR and CLEVR-Easy. Classic Sudokus have a 9x9 grid which is filled with digits from 1 to 9. In CLEVR-Sudoku these digits are replaced by images of objects. Hereby, a digit corresponds to a specific attribute combination, e.g., "yellow" and "sphere". Consequently, digits of the Sudoku are replaced by images of objects with these attribute combinations. These images each contain one object. To indicate, which attributes correspond to which digit, candidate examples of the digits are provided. The number of these examples is a flexible parameter, in our evaluations we used $N\in\{1,3,5,10\}$. Further, the number of images provided in the Sudoku grid is flexible as well. In our main evaluations we only considered CLEVR-Sudokus with $K=30$, meaning that 51 of the 81 Sudoku cells are filled and 30 are left to complete. For additional investigation we considered values for $K\in\{10,50\}$ as well. Examples of those Sudokus for Sudoku CLEVR are shown in Fig. 7. The dataset has a number of 1000 samples for Sudoku CLEVR-Easy and Sudoku CLEVR respectively for each value of $K$. Each sample has a different puzzle and a distinct set of images, no image is used twice for one puzzle [^5] The code for generating the dataset is available in our code repository, the already generated data files will be made public upon acceptance..
@@ E.3 Obtaining Revisory Feedback
  • Case 3: at least two clusters should not be deleted. In this case, we remove the encodings of these clusters completely. The cluster id for these clusters then no longer exist in the retrieval corpus.
@@ E.4 Neural Classification
Table 4: Classifying attributes from concept representations. Best results are in bold.
N Train SysBinder [[[[cont.]]]] SysBinder SysBinder (hard) SysBinder (step) NCB (P) NCB (P+E) NCB (P+E, topk)
@@
N=2000 $99.83\mbox{\scriptsize$\pm 0.24$}$ $92.49\mbox{\scriptsize$\pm 5.45$}$ $22.92\mbox{\scriptsize$\pm 0.00$}$ $95.76\mbox{\scriptsize$\pm 4.92$}$ $98.76\mbox{\scriptsize$\pm 1.05$}$ $\mathbf{99.02}\mbox{\scriptsize$\pm 1.00$}$ $98.93\mbox{\scriptsize$\pm 1.10$}$
N=200 $99.20\mbox{\scriptsize$\pm 0.41$}$ $87.90\mbox{\scriptsize$\pm 8.05$}$ $22.92\mbox{\scriptsize$\pm 0.00$}$ $92.42\mbox{\scriptsize$\pm 7.32$}$ $97.11\mbox{\scriptsize$\pm 2.16$}$ $\mathbf{98.50}\mbox{\scriptsize$\pm 1.80$}$ $98.42\mbox{\scriptsize$\pm 1.91$}$
N=50 $91.13\mbox{\scriptsize$\pm 4.21$}$ $78.41\mbox{\scriptsize$\pm 8.69$}$ $22.92\mbox{\scriptsize$\pm 0.00$}$ $70.64\mbox{\scriptsize$\pm 11.89$}$ $94.31\mbox{\scriptsize$\pm 4.47$}$ $\mathbf{95.87}\mbox{\scriptsize$\pm 2.93$}$ $95.72\mbox{\scriptsize$\pm 3.04$}$
N=20 $64.88\mbox{\scriptsize$\pm 10.89$}$ $62.61\mbox{\scriptsize$\pm 7.18$}$ $22.92\mbox{\scriptsize$\pm 0.00$}$ $54.61\mbox{\scriptsize$\pm 9.57$}$ $90.50\mbox{\scriptsize$\pm 7.09$}$ $\mathbf{94.22}\mbox{\scriptsize$\pm 4.11$}$ $94.15\mbox{\scriptsize$\pm 4.14$}$
@@
N=2000 $98.86\mbox{\scriptsize$\pm 1.15$}$ $86.22\mbox{\scriptsize$\pm 10.40$}$ $36.46\mbox{\scriptsize$\pm 0.00$}$ $88.90\mbox{\scriptsize$\pm 14.81$}$ $96.77\mbox{\scriptsize$\pm 2.63$}$ $\mathbf{97.26}\mbox{\scriptsize$\pm 2.67$}$ $97.17\mbox{\scriptsize$\pm 2.68$}$
N=200 $97.61\mbox{\scriptsize$\pm 2.58$}$ $81.13\mbox{\scriptsize$\pm 12.39$}$ $36.46\mbox{\scriptsize$\pm 0.00$}$ $83.17\mbox{\scriptsize$\pm 17.05$}$ $96.41\mbox{\scriptsize$\pm 2.64$}$ $\mathbf{96.80}\mbox{\scriptsize$\pm 3.01$}$ $96.80\mbox{\scriptsize$\pm 3.04$}$
N=50 $93.25\mbox{\scriptsize$\pm 4.62$}$ $61.67\mbox{\scriptsize$\pm 8.51$}$ $36.46\mbox{\scriptsize$\pm 0.00$}$ $68.81\mbox{\scriptsize$\pm 17.74$}$ $94.29\mbox{\scriptsize$\pm 4.78$}$ $\mathbf{94.67}\mbox{\scriptsize$\pm 4.65$}$ $94.10\mbox{\scriptsize$\pm 5.25$}$
N=20 $79.11\mbox{\scriptsize$\pm 8.75$}$ $49.79\mbox{\scriptsize$\pm 6.73$}$ $36.46\mbox{\scriptsize$\pm 0.00$}$ $58.58\mbox{\scriptsize$\pm 16.09$}$ $87.55\mbox{\scriptsize$\pm 5.35$}$ $\mathbf{88.57}\mbox{\scriptsize$\pm 4.68$}$ $88.42\mbox{\scriptsize$\pm 4.63$}$

### v2

@@ 4.1 Evaluations
Large, gray cube C4 $\land$ H5 $\land$ K5 $\land$ O13 $\land$ P6 (Gray1) $\land$ (Red $\lor$ Gray2) $\land$ (Large) $\land$ (Gray3) $\land$ (Gray4) “A large gray object”
Small, metal cube B4 $\land$ D4 $\land$ H1 $\land$ I1 $\land$ K1 (Cube) $\land$ (Small1) $\land$ (Small2) $\land$ (Small3) $\land$ (Small4) “A small cube
Large, blue sphere B1 $\land$ C7 $\land$ H4 $\land$ O1 $\land$ P2 (Sphere) $\land$ (Blue1) $\land$ (Blue2) $\land$ (Small $\lor$ Blue3) $\land$ (Blue4 $\lor$ Green $\lor$ Purple) “A blue sphere
Utilizing unsupervised concepts for understanding neural computations (Q4). In our final evaluations, we investigate whether NCB’s discrete concept encodings can make subsymbolic computations more transparent. We focus on the task of image classification using concept-bottleneck-like approaches [68, 34] on variations of the benchmark CLEVR-Hans3 dataset [68]. While the concept encodings in NCB + NN are trained unsupervised, they perform on par with the supervised approach of [68] (cf. Sec. F.7). More importantly, integrating NCB’s inherently inspectable concept representations into neural computations leads to more transparent decision processes.
@@
We illustrate this in Tab.  3, where we provide class-level explanations of the classifier in NCB + NN (cf. Sec. E.4 for details). Using NCB’s inspection mechanisms, human stakeholders can easily identify the classifier’s internal decision rules for a class (e.g., "a large gray object"). This is a critical feature for deploying trustworthy AI models in real-world scenarios. The key result is that this transparency is achieved even with unsupervised concept encodings. In Fig.  6, we further investigate whether a NCB-based neural classifier can be revised to mitigate confounders in the CLEVR-Hans3 dataset (cf. Sec. E.4 and Sec. F.8 for details). The confounding factor in the training set is the color gray, and we present the non-confounded test set accuracy in Fig.  6. We observe that standard loss-based feedback via explanatory interactive learning (XIL) [68] on the NN classifier’s explanations (+ XIL on NN) significantly reduces the effect of the confounder. Alternatively, by simply zeroing the activations of the undesired concept gray (+ XIL on concepts), we achieve even better confounding mitigation results without the typical issues of joint optimization. Our results highlight the potential of integrating NCB’s unsupervised concept representations for eliciting transparent and trustworthy subsymbolic computations. We thus answer Q4 affirmatively.
Limitations. NCB largely benefits from high-quality initial block-slot encodings. If these encodings are suboptimal, the resulting concept-slot encodings also degrade in quality. An important next step to handle more complex visual inputs, such as video data, is the integration of recent approaches (e.g., [16, 19]). Additionally, due to NCB’s unsupervised training nature, further alignment of NCB’s concepts is inevitable for effective deployment in downstream tasks [9]. Further, to build trust in NCB’s concept knowledge, human inspection is essential. Lastly, revisions are a critical aspect of NCB. However, they rely on humans to provide accurate feedback; a malicious user could manipulate NCB’s concepts. Fortunately, by inspecting the concept space, it is possible to track and mitigate such manipulation effectively.
@@ 5 Conclusions
In this work, we introduce the Neural Concept Binder framework for learning visual object-factor concepts in an unsupervised manner. Our evaluations suggest that NCB’s specific binding mechanisms facilitate the learning of expressive yet discrete concept representations. Furthermore, our results highlight the potential of integrating NCB’s inherently inspectable and revisable concept-slot encodings into both symbolic and neural modules. Promising directions for future research include exploring the benefits of NCB’s representations in continual learning settings [11], high-level concept learning [81], and probabilistic logic programming approaches [66, 67], as well as investigating connections to object-centric causal representation learning [48]. Lastly, incorporating downstream learning signals may be valuable (if present) for improving the quality of NCB’s initial concept encodings, e.g., through classification [5, 6] or differentiable clustering [76].
@@ Acknowledgments
The authors thank Gautam Singh for help with SysBinder and Cyprien Dzialo for preliminary results and insights. This work was supported by the Priority Program (SPP) 2422 in the subproject “Optimization of active surface design of high-speed progressive tools using machine and deep learning algorithms“ funded by the German Research Foundation (DFG), the ”ML2MT” project from the Volkswagen Stiftung and the ”The Adaptive Mind” project from the Hessian Ministry of Science and Arts (HMWK). It has further benefited from the HMWK projects ”The Third Wave of Artificial Intelligence - 3AI”, and Hessian.AI, as well as the Hessian research priority program LOEWE within the project WhiteBox, and the EU-funded “TANGO” project (EU Horizon 2023, GA No 57100431).
@@ Appendix A Details on Neural Concept Binder
A.1 Details on Systematic Binding and Slot Attention
The binding mechanism (SysBinder) of Singh et al. [65] allows images to be encoded into continuous block-slot representations and relies on the recently introduced slot attention mechanism [43]. In slot attention, so-called slots, $s\in R^{N_{S}\times N_{B}D_{B}}$ (each slot has dimension $N_{B}D_{B}$), compete for attending to parts of the input via a softmax-based attention. These slot encodings are iteratively updated and allow to capture distinct objects or image components. The result is an attention matrix $A\in R^{N_{S}\times D}$ for an input $x\in R^{D}$. Each entry $A_{i}$ corresponds to the attention weight of slot $i$ for the input $x$. Based on the attention matrix, the input is processed to read-out each object by multiplying $A$ with the input resulting in a matrix $U\in R^{N_{S}\times N_{B}D_{B}}$.
SysBinder now performs an additional factor binding on the vectors $u_{i}$ of $U$. The goal of this factor binding mechanism is to find a distribution over a codebook memory for each block in $u_{i}$, i.e., $u_{i}^{j}$. This codebook memory (one for each block), $M^{j}\in R^{K\times D_{B}}$, consists of a set of $K$ learnable codebook vectors. Specifically, for each block $j$ an RNN consisting of a GRU and MLP component iteratively updates the $j$-th block of slot $s_{i}$, $s_{i}^{j}$, based on $u_{i}^{j}$ and previous $s_{i}^{j}$. Finally, a soft information bottleneck is applied where each block $s_{i}^{j}$ performs dot-product attention over the codebook memory leading to the final block-slot representation:
[…]
This process is iteratively refined together with the refinement processes of slot attention. Overall, the encodings of SysBinder represent each object in an image by a slot with $N_{B}$ blocks where each block represents a factor of the object like shape or color.
Note that in the main text, the final $s_{i}^{j}$ is denoted as $z_{i}^{j}$.
A.2 Selection Function
@@ A.3 Details on Training
A.4 Instantiating Neural Concept Binder
We instantiate NCB’s soft binder via the SysBinder approach of  Singh et al. [65] which has been shown to provide valuable, object-factor disentangled representations. Thus, the soft binder was trained as in the original setup and with the published hyperparameters. Furthermore we instantiate the clustering model, $h$, via the powerful HDBSCAN method [12, 13, 54] (based on the popular HDBSCAN library [^2] https://hdbscan.readthedocs.io/en/latest/index.html). Hereby, $h$’s internal representation, $\phi$, consists of the learned hierarchical merge tree. In practice we found it beneficial to perform a grid search over $h$’s hyperparameters based on the unsupervised density-based cluster validity score [53]. The searched parameters are the minimal cluster size (the minimum number of samples in a group for that group to be considered a cluster) and minimal sample number (the number of samples in a neighborhood for a point to be considered as a core point) each over the values $[5,10,15,20,25,30,50,80,100]$. Moreover, we utilize the excess of mass algorithm and allow for single clusters. We performed the training of the retrieval corpus, i.e., fitting $h$, on a dataset of images containing single objects for simplifying the subsequent concept inspection mechanisms of our evaluations. However, this can easily be extended to multiple object images by utilising the soft binder’s slot attention masks to identify relevant objects in an image. Finally, we instantiate the retrieval corpus as a set of dictionaries and, unless stated otherwise, we utilise a retrieval corpus which contains one prototype and a set of exemplar encodings per concept. Furthermore, $s_{\mathcal{R}}$ represents the $\operatorname*{argmin}$ selection function and we utilize the euclidean distance as $d(\cdot,\cdot)$. It is important to note that $h$ does not make any assumptions about the number of clusters, $N_{C}$. Thus, although $h$ fits a clustering to best fit the block-slot encodings of a block, it can potentially provide an overparameterized clustering, e.g. by representing one underlying factor such as “gray” with several clusters. This highlights the importance of task-alignment, e.g., for symbolic downstream tasks, and concept inspection for general concept alignment. We refer to our code for more details [^3] Code available here., where trained model checkpoints and corresponding parameter logs are available.
A.5 Computational Resources
@@ Appendix B Details on CLEVR-Sudoku
CLEVR-Sudoku provides Sudokus based on the datasets CLEVR and CLEVR-Easy. Classic Sudokus have a 9x9 grid which is filled with digits from 1 to 9. In CLEVR-Sudoku these digits are replaced by images of objects. Hereby, a digit corresponds to a specific attribute combination, e.g., "yellow" and "sphere". Consequently, digits of the Sudoku are replaced by images of objects with these attribute combinations. These images each contain one object. To indicate, which attributes correspond to which digit, candidate examples of the digits are provided. The number of these examples is a flexible parameter, in our evaluations we used $N\in\{1,3,5,10\}$. Further, the number of images provided in the Sudoku grid is flexible as well. In our main evaluations we only considered CLEVR-Sudokus with $K=30$, meaning that 51 of the 81 Sudoku cells are filled and 30 are left to complete. For additional investigation we considered values for $K\in\{10,50\}$ as well. Examples of those Sudokus for Sudoku CLEVR are shown in Fig. 7. The dataset has a number of 1000 samples for Sudoku CLEVR-Easy and Sudoku CLEVR respectively for each value of $K$. Each sample has a different puzzle and a distinct set of images, no image is used twice for one puzzle [^4] The code for generating the dataset is available in our code repository, the already generated data files are accessible under https://huggingface.co/datasets/AIML-TUDA/CLEVR-Sudoku.
@@ Appendix D Baseline Models
NLOTM. NLOTM [80] builds on the principles of SysBinder and incorporates a Semantic Vector-Quantized (SVQ) Variational Autoencoder along with the Autoregressive LoT Prior (ALP). The SVQ component facilitates discrete semantic decomposition of a scene by learning hierarchical, composable factors that correspond closely to objects and their attributes in visual scenes. We refer to the original work for details.
@@ E.3 Obtaining Revisory Feedback
  • Case 3: at least two clusters should not be deleted. In this case, we completely remove the encodings of the to-be-removed clusters. The cluster id for these clusters no longer exists in the retrieval corpus.
@@ E.4 Neural Classification
Table 4: Ablation of NCB’s selection components for classifying attributes from concept representations. Best results are in bold.
N Train NCB (P) NCB (P+E) NCB (P+E, topk)
@@
N=2000 $98.76\mbox{\scriptsize$\pm 1.05$}$ $\mathbf{99.02}\mbox{\scriptsize$\pm 1.00$}$ $98.93\mbox{\scriptsize$\pm 1.10$}$
N=200 $97.11\mbox{\scriptsize$\pm 2.16$}$ $\mathbf{98.50}\mbox{\scriptsize$\pm 1.80$}$ $98.42\mbox{\scriptsize$\pm 1.91$}$
N=50 $94.31\mbox{\scriptsize$\pm 4.47$}$ $\mathbf{95.87}\mbox{\scriptsize$\pm 2.93$}$ $95.72\mbox{\scriptsize$\pm 3.04$}$
N=20 $90.50\mbox{\scriptsize$\pm 7.09$}$ $\mathbf{94.22}\mbox{\scriptsize$\pm 4.11$}$ $94.15\mbox{\scriptsize$\pm 4.14$}$
@@
N=2000 $96.77\mbox{\scriptsize$\pm 2.63$}$ $\mathbf{97.26}\mbox{\scriptsize$\pm 2.67$}$ $97.17\mbox{\scriptsize$\pm 2.68$}$
N=200 $96.41\mbox{\scriptsize$\pm 2.64$}$ $\mathbf{96.80}\mbox{\scriptsize$\pm 3.01$}$ $96.80\mbox{\scriptsize$\pm 3.04$}$
N=50 $94.29\mbox{\scriptsize$\pm 4.78$}$ $\mathbf{94.67}\mbox{\scriptsize$\pm 4.65$}$ $94.10\mbox{\scriptsize$\pm 5.25$}$
N=20 $87.55\mbox{\scriptsize$\pm 5.35$}$ $\mathbf{88.57}\mbox{\scriptsize$\pm 4.68$}$ $88.42\mbox{\scriptsize$\pm 4.63$}$